Circular No. 6472

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1991 A2 (ARAI) K. Dennerl, J. Englhauser, and J. Trümper, Max-Planck-Institut für

Extraterrestrische Physik, Garching, communicate: “X-ray emission from comet C/1991 A2 (= 1991b = 1990 XXVI) was detected in data obtained with the ROSAT Position Sensitive Proportional Counter during the allsky survey. The comet was observed from 1990 Nov. 18.09 to 21.43 UT (Δ = 1.2 AU, r = 1.5 AU), six weeks before it was discovered (IAUC 5157), when it would have had projected visual m_{1} ∼ 12 (or fainter) and would be optically the faintest comet ever detected in x-rays.”

γ-RAY BURST 960720 D. A. Frail, National Radio Astronomy Observatory; S. R. Kulkarni,

California Institute of Technology; L. Piro, E. Costa, M. Feroci, and M. Cinti, Istituto Astrofisica Spaziale, CNR, Frascati; F. Frontera, D. Dal Fiume, L. Nicastro, M. Orlandini, and G. Pizzichini, Istituto Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna; J. Heise and R. Jager, Space Research Organization of the Netherlands, Utrecht; and H. Muller, Beppo-SAX Scientific Data Center, Rome, report: “We made radio observations at the position of the γ-ray burst detected by Beppo-SAX on July 20 (IAUC 6467) using the Very Large Array (VLA) in Socorro at a wavelength of 20 cm. Images taken on Sept. 7.98 and 9.06 UT were compared with an image taken more than one year earlier as part of the VLA All-Sky Sur-

vey by Condon et al. (http://www.cv.nrao.edu/~jcondon/nvss.html).

There are no radio sources in the 10′ error circle above a 4σ level of 2 mJy in either the preburst or postburst images. This includes the position of the ROSAT x-ray source reported by Boller and Voges (IAUC 6469). There are five radio sources just outside the error circle but none shows any signs of unusual variability.”

1996 PW S. Mottola and U. Carsenty, DLR German Aerospace Establishment,

Berlin, write: “CCD photometry of 1996 PW, obtained during Sept. 7–10 with the Bochum 0.60-m telescope at the European Southern Observatory, reveals a rotational lightcurve with two maxima and two minima. We determined an unambiguous rotation period of 35.5 hr, an amplitude of 0.40 mag, and a reduced magnitude V (α = 8^{o}.5) = 14.69.”

1996 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green